OpenStudy (anonymous):

The price of a dozen eggs was $1.63. The price rose m dollars. Then the price dropped $.12 per dozen. Express the current price of eggs after the two price changes. A) m + 1.51 B) m + 1.63 C) m + 1.75 D) m + 2.83

OpenStudy (mathstudent55):

Price was 1.63 Price rose 0.12; newer price is 1.63 + m Price dropped $0.12; newest price is: 1.63 + m - 0.12 = 1.51 + m = m + 1.51